OpenStudy (anonymous):

simplify: (2-a^-1)^2

OpenStudy (anonymous):

It doesn't simplify nicely. (2-a^-1)² = (2 - 1/a)² = 4 - (4/a) + (1/a²) = (4a² - 4a + 1) / a²